i need to solve and graph the compound inequality if possible on a number line
-6<=3(x+5)<=18 please help

Answer by jim_thompson5910(35256)

-6 <= 3(x+5) <= 18


-6/3 <= 3(x+5)/3 <= 18/3 ... Divide all sides by 3


-2 <= x+5 <= 6


-2-5 <= x+5-5 <= 6-5 ... Subtract 5 from all sides


-7 <= x+0 <= 1


-7 <= x <= 1


-------------------------------------------------------


Graphing:


Step 1) Draw a number line. Include the values -10, -9, ..., 9, 10


Step 2) Plot closed circles at x = -7 and x = 1


Step 3) Shade between the closed circles.